Find the smallest whole number that when divided by 5,7,9, and 11, respectively, gives remainders of 1,2,3,and 4.
Answer by stanbon(48568)

Find the smallest whole number that when divided by 5,7,9, and 11, respectively, gives remainders of 1,2,3,and 4.
------------------------
Remainder of 1: 5*7*9*11 + 1 = 3465+1 = 3466
Remainder of 2: 3465+2 = 3467
Remainder of 3: 3465+3 = 3468
Remainder of 4: 3465+4 = 3469
==========================
Cheers,
Stan H.